OpenStudy (anonymous):

it is predicted that the population of a particular state will double every 25 years. a) determine the annual and monthly growth rate. express your answer as percents. b)determine the continuous growth rate per year. Express your answer as a percent.

OpenStudy (anonymous):

is this that half life thing?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

Find half life on google and it might explain more.

OpenStudy (anonymous):

If you start with 100 people, and you double every 25 yrs, then the formula is P = 100*2^(t/25) where the population is P at time t When t = 1, we get P = 100*2^(1/25) P = 102.81138 So we've gone from 100 to 102.81138, which means the growth rate is [(102.81138 - 100)/(100)] * 100 = 2.81138% So the annual growth rate is 2.81138%

OpenStudy (anonymous):

is the formula something like this:\[P(t)=P_0*2^(t/25)\]

OpenStudy (anonymous):

yeah thats what i was doing

OpenStudy (anonymous):

where do you learn this?

OpenStudy (anonymous):

i already knew this lol ;P

OpenStudy (anonymous):

im a smart lil girl lol

OpenStudy (anonymous):

yes you are haha

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

dont forget the pretty

OpenStudy (anonymous):

lmao thank you!!! :)

OpenStudy (anonymous):

you forgot to do the monthly growth rate? i got .231% what do you get?

OpenStudy (anonymous):

instead of t = 1, plug in t = 1/12 (since there are 12 months in a year) and simplify then compare that result with 100 to find the growth rate

OpenStudy (anonymous):

i knw that smart little girl :) i got .23%

OpenStudy (anonymous):

what about part "b" of the question?

OpenStudy (anonymous):

yes .23% is fine.... hold on im getting to part B lol

OpenStudy (anonymous):

i got 2.773% for part b but idk if I'm doing it right